Gil has $2 more than Karen. Together they have $70. How much money does each person have?
Answer by checkley77(12844)

x+x+2=70
2x=70-2
2x=68
x=68/2
x=$34 amount Karen had
34+2=$36 amount Gil had.
Proof:
34+36=70
70=70